OpenStudy (anonymous):

You purchase a lottery ticket for $2 from a grocery store. There are a total of 100,000 tickets in circulation. The grand prize is $10,000. What are your expected winnings? A. $0.10 B. $1.00 C. $2.00 D. $20

OpenStudy (anonymous):

Okauy, i know its not C.

OpenStudy (anonymous):

would my answer then be A?

OpenStudy (anonymous):

Is the money split between everyone who bought a ticket?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so would my expexted winnings be $1 then?

OpenStudy (anonymous):

@Jake-a-Shake

OpenStudy (anonymous):

@dumbcow would my answer be B?

OpenStudy (dumbcow):

expected value is sum of amount won * probability of winning \[EV = 9,998(\frac{1}{100,000}) -2(\frac{99,999}{100,000})\]

OpenStudy (anonymous):

do i solve that?

OpenStudy (dumbcow):

yeah just put it in calculator

OpenStudy (anonymous):

i dont have a calculator, gona use online one quick

OpenStudy (anonymous):

uggh stupid thibng won't let me solve the problem, it saus incomplete imput, but i put it in their just like how u had it

OpenStudy (dumbcow):

those answers are off or i missed something....you would expect to lose around $1.90 per ticket

OpenStudy (anonymous):

ya, unfortunetly i have to pick one of them:/

OpenStudy (anonymous):

its not 2.00 thats forsure, because I got that answer wrong last time so its between A, B, nd D

OpenStudy (dumbcow):

sure there isnt a typo or anything? im assuming only 1 ticket is picked for grand prize?

OpenStudy (anonymous):

yep im sure no typo.

OpenStudy (anonymous):

so if you were to pick one, which one you think it would be? 1.00?

OpenStudy (dumbcow):

i would not pick one because they are all wrong lol....you expect to lose , there are no expected winnings

OpenStudy (dumbcow):

go with A :)

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

Okay ill go with A then.

OpenStudy (dumbcow):

yes def A, i just saw my error i wasnt used to the wording, they dont want expected value i over complicated it expected winnings = amount won * probability of winning = 10,000*(1/100,000) = 1/10